The half-life of radioactive iodine 131 is approximately 8.3 days.
a) Find an exponential decay model for iodine 131.
b) following the fukushima Nuclear power plant disaster in 2011, radioactive iodine 131 was released into the atmosphere. How long would it have taken for 1500 grams of the iodine to decrease to 100 grams?

Found 2 solutions by josgarithmetic, Boreal:
Answer by josgarithmetic(39629)

This video will give the instruction needed:
Exponential Decay, example and instruction

Answer by Boreal(15235)

P=po*exp(-kt)
p= 1/2 po, because it is half as much
therefore, 1/2=exp(-kt)
ln both sides
ln(1/2)=-8.3k
0.693/8.3=k
do without rounding until end
0.0835=k
Here, you want (1/15) to be the ratio.
t is not known, but k is
therefore, (1/15)=exp(-0.0835*t)
ln both sides
t=32.43 days